# Starlink in Northern Europe: A New Look at Stationary and In-motion Performance

### v1

@@ Abstract
Index Terms:
ESIM, latency, mobility, satellite, throughputI Introduction
@@ III-B Software
This tool measures the maximum achievable throughput of Internet protocol (IP) networks. iperf3
supports measurements in both uplink and downlink directions, client-to-server and server-to-client, respectively. By default, iperf3
uses the CP, but also provides a UDP option. Additionally, iperf3
allows for the transmission of multiple parallel streams simultaneously, facilitating the assessment of the network’s maximum achievable performance. In this paper, we used TCP for both directions with the Windows default congestion control algorithm.
@@ III-B2 Ping
We used ping
to measure RTT. To be consistent with earlier studies [15, 9], in this paper, the term RTT is understood as two-way latency between the client and the server. For consistency, we used the same IP address
discussed in Section III-C3 for both ping
and iperf3
measurements. Additionally, we used the Windows default ping
setting, where a single ping
test sends uplink packets (each 32 bytes long) to a server and expects to receive the same number of packets as a reply from a server in the downlink.
@@
We used tracert
to understand how the uplink data travels from the client to the server. It helps to understand the following questions: (i) What is total number of hops between the client and server? (ii) Where is the nearest ground station? (iii) How does data travel from the ground station to the target server?
@@ III-C3 Servers locations
To investigate how Starlink data travels between the client and server and to assess the impact of different geographical locations on communications performance, we selected six iperf3
servers located at diverse geographical locations as shown in Table IV. Thanks to Hostkey666Hostkey B.V. Speedtest (https://speedtest.hostkey.com/), it offers iperf3
servers at five locations including (i) Helsinki, Finland; (ii) Amsterdam, Netherlands; (iii) Frankfurt, Germany; (iv) Reykjavík, Iceland; and (v) New York, United States. Notably, Starlink offers 100 % coverage to these cities777Starlink Availability Map (https://www.starlink.com/map).
@@
1 | Helsinki, Finland | spd-fisrv.hostkey.com | Hostkey |
2. | Amsterdam, Netherlands | spd-nlsrv.hostkey.com | Hostkey |
3. | Reykjavík, Iceland | spd-icsrv.hostkey.com | Hostkey |
4. | Frankfurt, Germany | spd-desrv.hostkey.com | Hostkey |
5. | New York, United States | spd-uswb.hostkey.com | Hostkey |
6. | Oulu, Finland | Anonymous | VTT |
Additionally, we set up a sixth iperf3
server in the same city as the client, Oulu, Finland which was running in a computer with Ubuntu 22.02
operating system. The computer was in VTT’s test network8885G and 6G test network environment by VTT (https://www.vttresearch.com/en/5g-and-6g-test-network-environment), and it was connected to the Internet using a 1 Gbps Ethernet interface. The computer has a public IP address and TCP port was open to the Internet during the test. The iperf3
version was 3.17 and we used a default configuration including TCP congestion control. Due to security reasons, we are not disclosing the IP address.
@@ IV-A Dataset and Service Plans
Our Starlink FHP terminal has 50 Gigabit/month Mobile Priority data subscription which was fully used during the measurements reported in Fig. 5. The analysis and measurements reported in Fig. 3, Fig. 7 and Fig. 8 were performed when the Mobile Priority data was already consumed. According to Starlink’s Fair Use Policy999Starlink Fair Use Policy (https://www.starlink.com/legal/documents/DOC-1469-65206-75), the user who has already consumed the allocated Mobile Priority data will be offered unlimited Mobile data for the rest of the month. For comparison, we also conducted tracert
and ping
tests using a 1 Gbps wired Internet connection.
@@ IV-B Stationary Measurements
In Fig. 3 ping
measurements consist of 774 samples for each server for Starlink. The terrestrial network measurements comprise the same number of samples. This gives us a total of 9,288 samples for both networks. In Fig. 5 Mobile Priority measurements were conducted at VTT, the Botanical Garden, and Auran Maja on May 24, 2024, while additional measurements at OuluZone were carried out on June 7, 2024. For each location and in each direction downlink/uplink, we collected 599 iperf3
measurement samples, resulting in a total of 4,792 measurement samples (averaged over ten iperf3
parallel streams) for all four locations.
For the results in Fig. 7, we conducted measurements on two consecutive days, June 26 and June 27, 2024. During the measurements, the vehicle was parked in the VTT parking lot, and we collected a total of 12,914 iperf3
measurement samples for uplink and downlink for all servers, except the one in New York. Specifically, we conducted an iperf3
test for each server, running separate uplink and downlink tests for 300 seconds. However, there were inconsistencies in the measurements sample length for each sever due to two reasons:
@@
From our five tests, we collected downlink iperf3
samples as 1,500, 1,500, 1,061, 1,212, and 1,200 for servers located in Oulu, Helsinki, Frankfurt, Amsterdam, and Reykjavík, respectively. The uplink dataset has 1,500, 1,222, 1,019, 1,500, and 1,200 samples for the same servers. This resulted in a total of 12,914 samples for the results presented in Fig. 7.
@@ IV-C In-motion measurements
For the in-motion measurements, we drove the vehicle five times on the path defined in Section III-C2. Four of these drive tests were conducted to collect iperf3
data, comprising two drive tests for uplink and two for downlink. The fifth drive test was performed to collect ping
samples. We collected 2068 in-motion samples, on average 830 iperf3
samples for uplink and downlink respectively, and 396 samples for ping
. During these drive tests, the average vehicle speed was around 50 km/h and the highest speed was 100 km/h.
@@ V-A Routing and Round-trip Time
Fig. 3 compares the number of hops (left) over the client to server route obtained from tracert
and ping
RTT (right) for terrestrial (top) and Starlink (bottom) networks from Oulu to six iperf3 servers as discussed in Section III-C3 and Table IV. When using the terrestrial network, there was only 1 hop and a relatively low RTT for the server in Oulu. This was because the client and server were installed on the same computer. When using Starlink, the server in Oulu experiences 18 hops. For the server in New York, both terrestrial and Starlink connection gives a comparable number of hops 19 and 20, respectively.
@@
Our tracert
analysis reveals that Starlink terminal in Oulu uses ground stations in Germany/Belgium. Therefore, the servers in Frankfurt and Amsterdam have a low number of hops compared to other locations. For all the servers including the one in New York, the traffic always passes through the IP address belonging to the USA as indicated in [12]. The data traffic started using the terrestrial network after being received by a ground station in Belgium or Germany. In this paper, we report Starlink tracert
results for a single iteration on July 24, 2024. Terrestrial tracert
results are for three iteration conducted on August 24, 2024. The number of routes may fluctuate slightly over time, potentially during the network congestion and the chosen service plan.
The right-hand side of the Fig. 3 illustrates boxchart101010Box chart (https://se.mathworks.com/help/matlab/ref/boxchart.html) representing the median RTT obtained from ping
measurements across these six locations. Each boxchart shows the following information (a) the median which is the line inside of each box, (a) the lower and upper quartiles represented by the top and bottom edges of each box (c) outliers illustrated using dots “.”, and (d) the minimum and maximum values excluding outliers indicated by the whiskers lines extend above and below each box. The ping
median results indicate that New York exhibits the highest median RTT of 103 ms and 143 ms for terrestrial and Starlink connections, respectively. This is followed by Raykjavik having median RTT of 75 ms and 99 ms for the terrestrial and Starlink connections, respectively. With Starlink connection, Frankfurt and Amsterdam demonstrate the lowest RTT with a median of 55 ms and 60 ms, respectively. This median RTT reduces to around 33 ms for the terrestrial network. Similarly, Oulu’s terrestrial median RTT is 0.028 ms which increases to 90 ms for Starlink because the data goes from Oulu to Frankfurt using satellite and returns from Frankfurt to Oulu using the terrestrial network. Compared to terrestrial connection, Starlink increases Helsinki’s median RTT from 15 ms to 85 ms. In Fig. 3 blue and red areas are references for Starlink’s expected performances in Standard and Mobile service plans, respectively. Except for the server in New York, the RTT for all other servers remained within RTT range supported by the Mobile service plan.
@@ V-B1 Stationary
Fig. 5 and Fig. 7 illustrate the Starlink uplink and downlink throughput performance. Specifically, Fig. 5 shows the Starlink downlink and uplink median throughput from four different test locations in Oulu to an iperf3
server in Finland. In the case of OuluZone, we use servers in Oulu and Helsinki. In Fig. 5 remaining measurements use only an iperf3
server in Helsinki. One can see that the measurement results are within the range of the Mobile Priority service plan. All test locations experienced median downlink and uplink throughput higher than 116 Mbps and 19.9 Mbps, respectively. The downlink performance for the Botanical Garden test location is significantly lower than the other test location due to possible LOS obstruction as indicated in Fig. 4. However, this trend could possibly be due to the terminal being connected to a Starlink satellite in an unobstructed area of FoV. The measurements at Auran Maja demonstrate the maximum downlink and uplink throughput 249 Mbps and 44 Mbps, respectively. Despite this, the median throughput remains closer to other other locations. Even though Mobile Priority data was already consumed, the measurement results in Fig. 7 align closely with the Mobile Priority service plan. This indicates the absence of interference and network congestion by other Starlink users during the measurements in Oulu. Moreover, these results reveal that Starlink connection is the bottleneck, and server location does not affect the throughput.

### v2

@@ Abstract
I Introduction
@@ III-B Software
This tool measures the maximum achievable throughput of Internet protocol (IP) networks. iperf3 supports measurements in both uplink and downlink directions, client-to-server and server-to-client, respectively. By default, iperf3 uses the CP, but also provides a UDP option. Additionally, iperf3 allows for the transmission of multiple parallel streams simultaneously, facilitating the assessment of the network’s maximum achievable performance. In this paper, we used TCP for both directions with the Windows default congestion control algorithm.
@@ III-B2 Ping
We used ping to measure RTT. To be consistent with earlier studies [15, 9], in this paper, the term RTT is understood as two-way latency between the client and the server. For consistency, we used the same IP address discussed in Section III-C3 for both ping and iperf3 measurements. Additionally, we used the Windows default ping setting, where a single ping test sends uplink packets (each 32 bytes long) to a server and expects to receive the same number of packets as a reply from a server in the downlink.
@@
We used tracert to understand how the uplink data travels from the client to the server. It helps to understand the following questions: (i) What is total number of hops between the client and server? (ii) Where is the nearest ground station? (iii) How does data travel from the ground station to the target server?
@@ III-C3 Servers locations
To investigate how Starlink data travels between the client and server and to assess the impact of different geographical locations on communications performance, we selected six iperf3 servers located at diverse geographical locations as shown in Table IV. Thanks to Hostkey666Hostkey B.V. Speedtest (https://speedtest.hostkey.com/), it offers iperf3 servers at five locations including (i) Helsinki, Finland; (ii) Amsterdam, Netherlands; (iii) Frankfurt, Germany; (iv) Reykjavík, Iceland; and (v) New York, United States. Notably, Starlink offers 100 % coverage to these cities777Starlink Availability Map (https://www.starlink.com/map).
@@
| 1 | Helsinki, Finland | spd-fisrv.hostkey.com | Hostkey |
| 2. | Amsterdam, Netherlands | spd-nlsrv.hostkey.com | Hostkey |
| 3. | Reykjavík, Iceland | spd-icsrv.hostkey.com | Hostkey |
| 4. | Frankfurt, Germany | spd-desrv.hostkey.com | Hostkey |
| 5. | New York, United States | spd-uswb.hostkey.com | Hostkey |
| 6. | Oulu, Finland | Anonymous | VTT |
Additionally, we set up a sixth iperf3 server in the same city as the client, Oulu, Finland which was running in a computer with Ubuntu 22.02 operating system. The computer was in VTT’s test network8885G and 6G test network environment by VTT (https://www.vttresearch.com/en/5g-and-6g-test-network-environment), and it was connected to the Internet using a 1 Gbps Ethernet interface. The computer has a public IP address and TCP port was open to the Internet during the test. The iperf3 version was 3.17 and we used a default configuration including TCP congestion control. Due to security reasons, we are not disclosing the IP address.
@@ IV-A Dataset and Service Plans
Our Starlink FHP terminal has 50 Gigabit/month Mobile Priority data subscription which was fully used during the measurements reported in Fig. 5. The analysis and measurements reported in Fig. 3, Fig. 7 and Fig. 8 were performed when the Mobile Priority data was already consumed. According to Starlink’s Fair Use Policy999Starlink Fair Use Policy (https://www.starlink.com/legal/documents/DOC-1469-65206-75), the user who has already consumed the allocated Mobile Priority data will be offered unlimited Mobile data for the rest of the month. For comparison, we also conducted tracert and ping tests using a 1 Gbps wired Internet connection.
@@ IV-B Stationary Measurements
In Fig. 3 ping measurements consist of 774 samples for each server for Starlink. The terrestrial network measurements comprise the same number of samples. This gives us a total of 9,288 samples for both networks. In Fig. 5 Mobile Priority measurements were conducted at VTT, the Botanical Garden, and Auran Maja on May 24, 2024, while additional measurements at OuluZone were carried out on June 7, 2024. For each location and in each direction downlink/uplink, we collected 599 iperf3 measurement samples, resulting in a total of 4,792 measurement samples (averaged over ten iperf3 parallel streams) for all four locations.
For the results in Fig. 7, we conducted measurements on two consecutive days, June 26 and June 27, 2024. During the measurements, the vehicle was parked in the VTT parking lot, and we collected a total of 12,914 iperf3 measurement samples for uplink and downlink for all servers, except the one in New York. Specifically, we conducted an iperf3 test for each server, running separate uplink and downlink tests for 300 seconds. However, there were inconsistencies in the measurements sample length for each sever due to two reasons:
@@
From our five tests, we collected downlink iperf3 samples as 1,500, 1,500, 1,061, 1,212, and 1,200 for servers located in Oulu, Helsinki, Frankfurt, Amsterdam, and Reykjavík, respectively. The uplink dataset has 1,500, 1,222, 1,019, 1,500, and 1,200 samples for the same servers. This resulted in a total of 12,914 samples for the results presented in Fig. 7.
@@ IV-C In-motion measurements
For the in-motion measurements, we drove the vehicle five times on the path defined in Section III-C2. Four of these drive tests were conducted to collect iperf3 data, comprising two drive tests for uplink and two for downlink. The fifth drive test was performed to collect ping samples. We collected 2068 in-motion samples, on average 830 iperf3 samples for uplink and downlink respectively, and 396 samples for ping. During these drive tests, the average vehicle speed was around 50 km/h and the highest speed was 100 km/h.
@@ V-A Routing and Round-trip Time
Fig. 3 compares the number of hops (left) over the client to server route obtained from tracert and ping RTT (right) for terrestrial (top) and Starlink (bottom) networks from Oulu to six iperf3 servers as discussed in Section III-C3 and Table IV. When using the terrestrial network, there was only 1 hop and a relatively low RTT for the server in Oulu. This was because the client and server were installed on the same computer. When using Starlink, the server in Oulu experiences 18 hops. For the server in New York, both terrestrial and Starlink connection gives a comparable number of hops 19 and 20, respectively.
@@
Our tracert analysis reveals that Starlink terminal in Oulu uses ground stations in Germany/Belgium. Therefore, the servers in Frankfurt and Amsterdam have a low number of hops compared to other locations. For all the servers including the one in New York, the traffic always passes through the IP address belonging to the USA as indicated in [12]. The data traffic started using the terrestrial network after being received by a ground station in Belgium or Germany. In this paper, we report Starlink tracert results for a single iteration on July 24, 2024. Terrestrial tracert results are for three iteration conducted on August 24, 2024. The number of routes may fluctuate slightly over time, potentially during the network congestion and the chosen service plan.
The right-hand side of the Fig. 3 illustrates boxchart101010Box chart (https://se.mathworks.com/help/matlab/ref/boxchart.html) representing the median RTT obtained from ping measurements across these six locations. Each boxchart shows the following information (a) the median which is the line inside of each box, (a) the lower and upper quartiles represented by the top and bottom edges of each box (c) outliers illustrated using dots “.”, and (d) the minimum and maximum values excluding outliers indicated by the whiskers lines extend above and below each box. The ping median results indicate that New York exhibits the highest median RTT of 103 ms and 143 ms for terrestrial and Starlink connections, respectively. This is followed by Raykjavik having median RTT of 75 ms and 99 ms for the terrestrial and Starlink connections, respectively. With Starlink connection, Frankfurt and Amsterdam demonstrate the lowest RTT with a median of 55 ms and 60 ms, respectively. This median RTT reduces to around 33 ms for the terrestrial network. Similarly, Oulu’s terrestrial median RTT is 0.028 ms which increases to 90 ms for Starlink because the data goes from Oulu to Frankfurt using satellite and returns from Frankfurt to Oulu using the terrestrial network. Compared to terrestrial connection, Starlink increases Helsinki’s median RTT from 15 ms to 85 ms. In Fig. 3 blue and red areas are references for Starlink’s expected performances in Standard and Mobile service plans, respectively. Except for the server in New York, the RTT for all other servers remained within RTT range supported by the Mobile service plan.
@@ V-B1 Stationary
Fig. 5 and Fig. 7 illustrate the Starlink uplink and downlink throughput performance. Specifically, Fig. 5 shows the Starlink downlink and uplink median throughput from four different test locations in Oulu to an iperf3 server in Finland. In the case of OuluZone, we use servers in Oulu and Helsinki. In Fig. 5 remaining measurements use only an iperf3 server in Helsinki. One can see that the measurement results are within the range of the Mobile Priority service plan. All test locations experienced median downlink and uplink throughput higher than 116 Mbps and 19.9 Mbps, respectively. The downlink performance for the Botanical Garden test location is significantly lower than the other test location due to possible LOS obstruction as indicated in Fig. 4. However, this trend could possibly be due to the terminal being connected to a Starlink satellite in an unobstructed area of FoV. The measurements at Auran Maja demonstrate the maximum downlink and uplink throughput 249 Mbps and 44 Mbps, respectively. Despite this, the median throughput remains closer to other other locations. Even though Mobile Priority data was already consumed, the measurement results in Fig. 7 align closely with the Mobile Priority service plan. This indicates the absence of interference and network congestion by other Starlink users during the measurements in Oulu. Moreover, these results reveal that Starlink connection is the bottleneck, and server location does not affect the throughput.